\begin{thebibliography}{10}

\bibitem{Sofue2001}
Y.~Sofue and V.~Rubin, ``{Rotation curves of spiral galaxies},'' {\em Annual
  Review of Astronomy and Astrophysics}, vol.~39, pp.~137--174, sep 2001.

\bibitem{1978AJ.....83.1050S}
G.~Steigman, H.~Quintana, C.~L. Sarazin, and J.~Faulkner, ``{Dynamical
  interactions and astrophysical effects of stable heavy neutrinos},'' {\em The
  Astronomical Journal}, vol.~83, p.~1050, sep 1978.

\bibitem{Read_2018fxs}
J.~I. Read, M.~G. Walker, and P.~Steger, ``{Dark matter heats up in dwarf
  galaxies},'' {\em Monthly Notices of the Royal Astronomical Society},
  vol.~484, no.~1, pp.~1401--1420, 2019.

\bibitem{Hassani_2020zvz}
E.~Hassani, R.~Pazhouhesh, and H.~Ebadi, ``{The effect of dark matter on stars
  at the Galactic center: The paradox of youth problem},'' {\em International
  Journal of Modern Physics D}, vol.~29, p.~2050052, jun 2020.

\bibitem{2020PhRvD.102b3024C}
M.~H. Chan and C.~M. Lee, ``{Constraining the spin-independent elastic
  scattering cross section of dark matter using the Moon as a detection target
  and the background neutrino data},'' {\em Physical Review D}, vol.~102,
  no.~2, p.~23024, 2020.

\bibitem{2020PhLB..80435403G}
R.~Garani and P.~Tinyakov, ``{Constraints on dark matter from the Moon},'' {\em
  Physics Letters, Section B: Nuclear, Elementary Particle and High-Energy
  Physics}, vol.~804, p.~135403, may 2020.

\bibitem{2012JCAP...07..046H}
D.~Hooper and J.~H. Steffen, ``{Dark matter and the habitability of planets},''
  {\em Journal of Cosmology and Astroparticle Physics}, vol.~2012, no.~7,
  p.~46, 2012.

\bibitem{2012PhRvD..85b3519M}
S.~D. McDermott, H.~B. Yu, and K.~M. Zurek, ``{Constraints on scalar asymmetric
  dark matter from black hole formation in neutron stars},'' {\em Physical
  Review D - Particles, Fields, Gravitation and Cosmology}, vol.~85, p.~23519,
  jan 2012.

\bibitem{2011IJMPE..20..109R}
M.~RAZEIRA, A.~MESQUITA, C.~A.~Z. VASCONCELLOS, and R.~O. GOMES, ``{Accretion
  of Dark Matter in Neutron Stars},'' {\em International Journal of Modern
  Physics E}, vol.~20, no.~supp02, pp.~109--116, 2011.

\bibitem{2010PhRvL.105n1101P}
M.~A. Perez-Garcia, J.~Silk, and J.~R. Stone, ``{Dark matter, neutron stars,
  and strange quark matter},'' {\em Physical Review Letters}, vol.~105, no.~14,
  p.~141101, 2010.

\bibitem{2010PhRvD..81l3521D}
A.~{De Lavallaz} and M.~Fairbairn, ``{Neutron stars as dark matter probes},''
  {\em Physical Review D - Particles, Fields, Gravitation and Cosmology},
  vol.~81, p.~123521, jun 2010.

\bibitem{Bertone_2008}
G.~Bertone and M.~Fairbairn, ``{Compact stars as dark matter probes},'' {\em
  Physical Review D - Particles, Fields, Gravitation and Cosmology}, vol.~77,
  p.~43515, feb 2008.

\bibitem{2008PhRvD..77b3006K}
C.~Kouvaris, ``{WIMP annihilation and cooling of neutron stars},'' {\em
  Physical Review D - Particles, Fields, Gravitation and Cosmology}, vol.~77,
  p.~23006, jan 2008.

\bibitem{Bell:2020jou}
N.~F. Bell, G.~Busoni, S.~Robles, and M.~Virgato, ``{Improved treatment of dark
  matter capture in neutron stars},'' {\em Journal of Cosmology and
  Astroparticle Physics}, vol.~2020, no.~09, pp.~028--028, 2020.

\bibitem{Joglekar:2020liw}
A.~Joglekar, N.~Raj, P.~Tanedo, and H.-B. Yu, ``{Kinetic Heating from Contact
  Interactions with Relativistic Targets: Electrons Capture Dark Matter in
  Neutron Stars},'' 2020.

\bibitem{2020arXiv200405312L}
G.~L. Lin and Y.~H. Lin, ``{Analysis on the black hole formations inside old
  neutron stars by isospin-violating dark matter with self-interaction},'' {\em
  Journal of Cosmology and Astroparticle Physics}, vol.~2020, no.~8,
  pp.~022--022, 2020.

\bibitem{2019PhRvD.100j3019H}
A.~Herrero, M.~A. P{\'{e}}rez-Garc{\'{i}}a, J.~Silk, and C.~Albertus, ``{Dark
  matter and bubble nucleation in old neutron stars},'' {\em Physical Review
  D}, vol.~100, no.~10, p.~103019, 2019.

\bibitem{2019JCAP...06..054B}
N.~F. Bell, G.~Busoni, and S.~Robles, ``{Capture of leptophilic dark matter in
  neutron stars},'' {\em Journal of Cosmology and Astroparticle Physics},
  vol.~2019, p.~54, jun 2019.

\bibitem{2020JCAP...06..007F}
M.~Fabbrichesi and A.~Urbano, ``{Charged neutron stars and observational tests
  of a dark force weaker than gravity},'' {\em Journal of Cosmology and
  Astroparticle Physics}, vol.~2020, no.~6, p.~7, 2020.

\bibitem{2018JCAP...09..018B}
N.~F. Bell, G.~Busoni, and S.~Robles, ``{Heating up neutron stars with
  inelastic dark matter},'' {\em Journal of Cosmology and Astroparticle
  Physics}, vol.~2018, no.~9, p.~18, 2018.

\bibitem{2018JHEP...11..096K}
J.~Kopp, R.~Laha, T.~Opferkuch, and W.~Shepherd, ``{Cuckoo's eggs in neutron
  stars: can LIGO hear chirps from the dark sector?},'' {\em Journal of High
  Energy Physics}, vol.~2018, no.~11, p.~96, 2018.

\bibitem{2017PASA...34...43C}
M.~Cerme{\~{n}}o, M.~{\'{A}}. P{\'{e}}rez-Garc{\'{i}}a, and J.~Silk,
  ``{Fermionic Light Dark Matter Particles and the New Physics of Neutron
  Stars},'' {\em Publications of the Astronomical Society of Australia},
  vol.~34, p.~e043, 2017.

\bibitem{2017arXiv170508864H}
X.~Huang, J.-F. Liu, W.-H. Wang, Q.~Cheng, and X.-P. Zheng, ``{Bondi accretion
  of dark matter by neutron stars},'' {\em arXiv e-prints},
  p.~arXiv:1705.08864, 2017.

\bibitem{2018ApJ...863..157C}
M.~Cerme{\~{n}}o, M.~A. P{\'{e}}rez-Garc{\'{i}}a, and R.~A. Lineros,
  ``{Enhanced Neutrino Emissivities in Pseudoscalar-mediated Dark Matter
  Annihilation in Neutron Stars},'' {\em The Astrophysical Journal}, vol.~863,
  no.~2, p.~157, 2018.

\bibitem{2017PhRvL.119m1801B}
M.~Baryakhtar, J.~Bramante, S.~W. Li, T.~Linden, and N.~Raj, ``{Dark Kinetic
  Heating of Neutron Stars and an Infrared Window on WIMPs, SIMPs, and Pure
  Higgsinos},'' {\em Physical Review Letters}, vol.~119, no.~13, p.~131801,
  2017.

\bibitem{Perez-Garcia:2015rta}
M.~A. Perez-Garcia, ``{Inhomogeneous seeding of quark bubbles in Neutron
  Stars},'' in {\em Compact Stars in the QCD Phase Diagram IV}, 2015.

\bibitem{Perez-Garcia:2014nza}
M.~P{\'{e}}rez-Garc{\'{i}}a and J.~Silk, ``{Dark matter annihilation and local
  warming in the core of a neutron star},'' in {\em 20th International
  Conference on Particles and Nuclei}, pp.~365--368, 2014.

\bibitem{2015ApJ...800..141Z}
H.~Zheng, K.~J. Sun, and L.~W. Chen, ``{Old neutron stars as probes of
  isospin-violating dark matter},'' {\em Astrophysical Journal}, vol.~800,
  no.~2, p.~141, 2015.

\bibitem{Tachibana:2013iva}
M.~Tachibana and M.~Ruggieri, ``{Comment on Dark Matter Capture in Neutron
  Stars with Exotic Phases},'' in {\em arXiv:1312.5802 [hep-ph]}, 2013.

\bibitem{2013arXiv1306.0148P}
M.~A. Perez-Garcia, ``{Deconfinement of quark content in neutron stars due to
  cosmic external agents},'' {\em arXiv e-prints}, p.~arXiv:1306.0148, 2013.

\bibitem{Perez-Garcia:2013dwa}
M.~{\'{A}}. P{\'{e}}rez-Garc{\'{i}}a, F.~Daigne, and J.~Silk, ``{Short
  gamma-ray bursts and dark matter seeding in neutron stars},'' {\em
  Astrophysical Journal}, vol.~768, no.~2, p.~145, 2013.

\bibitem{2013PhRvD..87l3507B}
N.~F. Bell, A.~Melatos, and K.~Petraki, ``{Realistic neutron star constraints
  on bosonic asymmetric dark matter},'' {\em Physical Review D - Particles,
  Fields, Gravitation and Cosmology}, vol.~87, p.~123507, jun 2013.

\bibitem{2013PhRvD..87e5012B}
J.~Bramante, K.~Fukushima, and J.~Kumar, ``{Constraints on bosonic dark matter
  from observation of old neutron stars},'' {\em Physical Review D - Particles,
  Fields, Gravitation and Cosmology}, vol.~87, p.~55012, mar 2013.

\bibitem{Serebrov:2012vm}
A.~P. Serebrov, O.~M. Zherebtsov, A.~K. Fomin, and M.~S. Onegin,
  ``{Ultracold-neutron trap as a tool to search for dark matter interacting
  with long-range forces},'' {\em Physical Review C - Nuclear Physics},
  vol.~86, no.~2, p.~25802, 2012.

\bibitem{2012AIPC.1441..525P}
M.~{\'{A}}. P{\'{e}}rez-Garc{\'{i}}a, J.~Silk, and J.~R. Stone, ``{Dark matter
  seeding in neutron stars},'' in {\em AIP Conference Proceedings} (S.~G.
  Steadman, G.~S. Stephans, and F.~E. Taylor, eds.), vol.~1441 of {\em American
  Institute of Physics Conference Series}, pp.~525--527, 2012.

\bibitem{Wong:2011ege}
K.~W. Wong, {\em {Physical Properties of White Dwarf with a Dark Matter Core}}.
\newblock PhD thesis, Hong Kong, Chinese U., 2011.

\bibitem{2019JCAP...08..018D}
B.~Dasgupta, A.~Gupta, and A.~Ray, ``{Dark matter capture in celestial objects:
  improved treatment of multiple scattering and updated constraints from white
  dwarfs},'' {\em Journal of Cosmology and Astroparticle Physics}, vol.~2019,
  pp.~018--018, aug 2019.

\bibitem{2018PhRvD..98f3002C}
M.~Cerme{\~{n}}o and M.~A. P{\'{e}}rez-Garc{\'{i}}a, ``{Gamma rays from dark
  mediators in white dwarfs},'' {\em Physical Review D}, vol.~98, no.~6,
  p.~63002, 2018.

\bibitem{2018PhRvD..98k5027G}
P.~W. Graham, R.~Janish, V.~Narayan, S.~Rajendran, and P.~Riggins, ``{White
  dwarfs as dark matter detectors},'' {\em Physical Review D}, vol.~98, no.~11,
  p.~115027, 2018.

\bibitem{2018PhRvD..98j3023N}
J.~S. Niu, T.~Li, W.~Zong, H.~F. Xue, and Y.~Wang, ``{Probing the dark
  matter-electron interactions via hydrogen-atmosphere pulsating white
  dwarfs},'' {\em Physical Review D}, vol.~98, no.~10, p.~103023, 2018.

\bibitem{2017arXiv170310104N}
J.-S. Niu, W.~Zong, and H.-F. Xue, ``{Constrain the Dark Matter Electron Cross
  Section from Pulsating White Dwarfs},'' {\em arXiv e-prints},
  p.~arXiv:1703.10104, 2017.

\bibitem{2016MNRAS.459..695A}
P.~Amaro-Seoane, J.~Casanellas, R.~Sch{\"{o}}del, E.~Davidson, and J.~Cuadra,
  ``{Probing dark matter crests with white dwarfs and IMBHs},'' {\em Monthly
  Notices of the Royal Astronomical Society}, vol.~459, no.~1, pp.~695--700,
  2016.

\bibitem{2015PhRvD..91j3514H}
T.~J. Hurst, A.~R. Zentner, A.~Natarajan, and C.~Badenes, ``{Indirect probes of
  dark matter and globular cluster properties from dark matter annihilation
  within the coolest white dwarfs},'' {\em Physical Review D - Particles,
  Fields, Gravitation and Cosmology}, vol.~91, no.~10, p.~103514, 2015.

\bibitem{2013PhRvD..87l3506L}
S.~C. Leung, M.~C. Chu, L.~M. Lin, and K.~W. Wong, ``{Dark-matter admixed white
  dwarfs},'' {\em Physical Review D - Particles, Fields, Gravitation and
  Cosmology}, vol.~87, p.~123506, jun 2013.

\bibitem{2009JCAP...08..024U}
H.~Umeda, N.~Yoshida, K.~Nomoto, S.~Tsuruta, M.~Sasaki, and T.~Ohkubo, ``{Early
  black hole formation by accretion of gas and dark matter},'' {\em Journal of
  Cosmology and Astroparticle Physics}, vol.~2009, no.~8, p.~24, 2009.

\bibitem{Belotsky2014}
K.~M. Belotsky, A.~E. Dmitriev, E.~A. Esipova, V.~A. Gani, A.~V. Grobov, M.~Y.
  Khlopov, A.~A. Kirillov, S.~G. Rubin, and I.~V. Svadkovsky, ``{Signatures of
  primordial black hole dark matter},'' {\em Modern Physics Letters A},
  vol.~29, p.~1440005, dec 2014.

\bibitem{Hassani2020}
E.~Hassani, H.~Ebadi, and R.~Pazhouhesh, ``{Capture rate of weakly interacting
  massive particles (WIMPs) in binary star systems},'' sep 2020.

\bibitem{1985ApJ...296..679P}
W.~H. Press and D.~N. Spergel, ``{Capture by the sun of a galactic population
  of weakly interacting, massive particles},'' {\em The Astrophysical Journal},
  vol.~296, p.~679, sep 1985.

\bibitem{Gould_1987ir}
A.~Gould, ``{Resonant enhancements in weakly interacting massive particle
  capture by the earth},'' {\em The Astrophysical Journal}, vol.~321, p.~571,
  oct 1987.

\bibitem{PhysRevD.96.063001}
L.~G. G{\'{o}}mez and J.~A. Rueda, ``{Dark matter dynamical friction versus
  gravitational wave emission in the evolution of compact-star binaries},''
  {\em Physical Review D}, vol.~96, no.~6, p.~63001, 2017.

\bibitem{Penarrubia2016}
J.~Pe{\~{n}}arrubia, A.~D. Ludlow, J.~Chanam{\'{e}}, and M.~G. Walker, ``{Wide
  binaries in ultrafaint galaxies: A window on to dark matter on the smallest
  scales},'' {\em Monthly Notices of the Royal Astronomical Society: Letters},
  vol.~461, pp.~L72--L76, may 2016.

\bibitem{Blas2020}
D.~Blas, D.~L. Nacir, and S.~Sibiryakov, ``{Secular effects of ultralight dark
  matter on binary pulsars},'' {\em Physical Review D}, vol.~101, p.~063016,
  mar 2020.

\bibitem{GOMEZ2019100343}
L.~G. G{\'{o}}mez, ``{Constraining light fermionic dark matter with binary
  pulsars},'' {\em Physics of the Dark Universe}, vol.~26, p.~100343, 2019.

\bibitem{Blas2017}
D.~Blas, D.~L. Nacir, and S.~Sibiryakov, ``{Ultralight Dark Matter Resonates
  with Binary Pulsars},'' {\em Physical Review Letters}, vol.~118, jun 2017.

\bibitem{2015PhRvD..92l3530P}
P.~Pani, ``{Binary pulsars as dark-matter probes},'' {\em Physical Review D -
  Particles, Fields, Gravitation and Cosmology}, vol.~92, no.~12, p.~123530,
  2015.

\bibitem{Armaleo2020}
J.~M. Armaleo, D.~L. Nacir, and F.~R. Urban, ``{Binary pulsars as probes for
  spin-2 ultralight dark matter},'' {\em Journal of Cosmology and Astroparticle
  Physics}, vol.~2020, p.~053, jan 2020.

\bibitem{CAPUTO20181}
A.~Caputo, J.~Zavala, and D.~Blas, ``{Binary pulsars as probes of a Galactic
  dark matter disk},'' {\em Physics of the Dark Universe}, vol.~19, pp.~1--11,
  mar 2018.

\bibitem{Yoo2004}
J.~Yoo, J.~Chaname, and A.~Gould, ``{The End of the MACHO Era: Limits on Halo
  Dark Matter from Stellar Halo Wide Binaries},'' {\em The Astrophysical
  Journal}, vol.~601, pp.~311--318, jan 2004.

\bibitem{1996PhR...267..195J}
G.~Jungman, M.~Kamionkowski, and K.~Griest, ``{Supersymmetric dark matter},''
  {\em Physics Report}, vol.~267, pp.~195--373, mar 1996.

\bibitem{Lopes2011}
I.~Lopes, J.~Casanellas, and D.~Eug{\'{e}}nio, ``{The capture of dark matter
  particles through the evolution of low-mass stars},'' {\em Physical Review D
  - Particles, Fields, Gravitation and Cosmology}, vol.~83, p.~63521, mar 2011.

\bibitem{Bond2017}
H.~E. Bond, G.~H. Schaefer, R.~L. Gilliland, J.~B. Holberg, B.~D. Mason, I.~W.
  Lindenblad, M.~Seitz-McLeese, W.~D. Arnett, P.~Demarque, F.~Spada, P.~A.
  Young, M.~A. Barstow, M.~R. Burleigh, and D.~Gudehus, ``{ The Sirius System
  and Its Astrophysical Puzzles: Hubble Space Telescope and Ground-based
  Astrometry },'' {\em The Astrophysical Journal}, vol.~840, p.~70, may 2017.

\bibitem{Holberg1998}
J.~B. Holberg, M.~A. Barstow, F.~C. Bruhweiler, A.~M. Cruise, and A.~J. Penny,
  ``{Sirius B: A New, More Accurate View},'' {\em The Astrophysical Journal},
  vol.~497, pp.~935--942, apr 1998.

\bibitem{2014JCAP...05..013G}
T.~G{\"{u}}ver, A.~E. Erkoca, M.~H. Reno, and I.~Sarcevic, ``{On the capture of
  dark matter by neutron stars},'' {\em Journal of Cosmology and Astroparticle
  Physics}, vol.~2014, pp.~013--013, jan 2014.

\bibitem{2011PhRvL.106b1303B}
E.~Behnke, J.~Behnke, S.~J. Brice, D.~Broemmelsiek, J.~I. Collar, P.~S. Cooper,
  M.~Crisler, C.~E. Dahl, D.~Fustin, J.~Hall, J.~H. Hinnefeld, M.~Hu,
  I.~Levine, E.~Ramberg, T.~Shepherd, A.~Sonnenschein, and M.~Szydagis,
  ``{Improved limits on spin-dependent WIMP-proton interactions from a two
  liter CF3I bubble chamber},'' {\em Physical Review Letters}, vol.~106,
  p.~21303, jan 2011.

\bibitem{hilditch_2001}
R.~W. Hilditch, {\em {An Introduction to Close Binary Stars}}.
\newblock Cambridge University Press, mar 2001.

\bibitem{Freire2011}
P.~C. Freire, C.~G. Bassa, N.~Wex, I.~H. Stairs, D.~J. Champion, S.~M. Ransom,
  P.~Lazarus, V.~M. Kaspi, J.~W. Hessels, M.~Kramer, J.~M. Cordes, J.~P.
  Verbiest, P.~Podsiadlowski, D.~J. Nice, J.~S. Deneva, D.~R. Lorimer, B.~W.
  Stappers, M.~A. Mclaughlin, and F.~Camilo, ``{On the nature and evolution of
  the unique binary pulsar J1903+0327},'' {\em Monthly Notices of the Royal
  Astronomical Society}, vol.~412, pp.~2763--2780, apr 2011.

\bibitem{Wex2014}
N.~Wex, ``{Testing relativistic gravity with radio pulsars},'' in {\em
  Applications and Experiments}, DE GRUYTER, aug 2014.

\bibitem{Arzoumanian2018}
Z.~Arzoumanian, A.~Brazier, S.~Burke-Spolaor, S.~Chamberlin, S.~Chatterjee,
  B.~Christy, J.~M. Cordes, N.~J. Cornish, F.~Crawford, H.~T. Cromartie,
  K.~Crowter, M.~E. DeCesar, P.~B. Demorest, T.~Dolch, J.~A. Ellis, R.~D.
  Ferdman, E.~C. Ferrara, E.~Fonseca, N.~Garver-Daniels, P.~A. Gentile,
  D.~Halmrast, E.~A. Huerta, F.~A. Jenet, C.~Jessup, G.~Jones, M.~L. Jones,
  D.~L. Kaplan, M.~T. Lam, T.~J. {W. Lazio}, L.~Levin, A.~Lommen, D.~R.
  Lorimer, J.~Luo, R.~S. Lynch, D.~Madison, A.~M. Matthews, M.~A. McLaughlin,
  S.~T. McWilliams, C.~Mingarelli, C.~Ng, D.~J. Nice, T.~T. Pennucci, S.~M.
  Ransom, P.~S. Ray, X.~Siemens, J.~Simon, R.~Spiewak, I.~H. Stairs, D.~R.
  Stinebring, K.~Stovall, J.~K. Swiggum, S.~R. Taylor, M.~Vallisneri, R.~van
  Haasteren, S.~J. Vigeland, and W.~Zhu, ``{The NANOGrav 11-year Data Set:
  High-precision Timing of 45 Millisecond Pulsars},'' {\em The Astrophysical
  Journal Supplement Series}, vol.~235, p.~37, apr 2018.

\bibitem{Desvignes2016}
G.~Desvignes, R.~N. Caballero, L.~Lentati, J.~P. Verbiest, D.~J. Champion,
  B.~W. Stappers, G.~H. Janssen, P.~Lazarus, S.~Os{\l}owski, S.~Babak, C.~G.
  Bassa, P.~Brem, M.~Burgay, I.~Cognard, J.~R. Gair, E.~Graikou, L.~Guillemot,
  J.~W. Hessels, A.~Jessner, C.~Jordan, R.~Karuppusamy, M.~Kramer, A.~Lassus,
  K.~Lazaridis, K.~J. Lee, K.~Liu, A.~G. Lyne, J.~McKee, C.~M. Mingarelli,
  D.~Perrodin, A.~Petiteau, A.~Possenti, M.~B. Purver, P.~A. Rosado,
  S.~Sanidas, A.~Sesana, G.~Shaifullah, R.~Smits, S.~R. Taylor, G.~Theureau,
  C.~Tiburzi, R.~{Van Haasteren}, and A.~Vecchio, ``{High-precision timing of
  42 millisecond pulsars with the European Pulsar Timing Array},'' {\em Monthly
  Notices of the Royal Astronomical Society}, vol.~458, pp.~3341--3380, mar
  2016.

\bibitem{Gonzalez2011}
M.~E. Gonzalez, I.~H. Stairs, R.~D. Ferdman, P.~C. Freire, D.~J. Nice, P.~B.
  Demorest, S.~M. Ransom, M.~Kramer, F.~Camilo, G.~Hobbs, R.~N. Manchester, and
  A.~G. Lyne, ``{High-precision timing of five millisecond pulsars: Space
  velocities, binary evolution, and equivalence principles},'' {\em
  Astrophysical Journal}, vol.~743, dec 2011.

\bibitem{2013Sci...340..448A}
J.~Antoniadis, P.~C. Freire, N.~Wex, T.~M. Tauris, R.~S. Lynch, M.~H. {Van
  Kerkwijk}, M.~Kramer, C.~Bassa, V.~S. Dhillon, T.~Driebe, J.~W. Hessels,
  V.~M. Kaspi, V.~I. Kondratiev, N.~Langer, T.~R. Marsh, M.~A. McLaughlin,
  T.~T. Pennucci, S.~M. Ransom, I.~H. Stairs, J.~{Van Leeuwen}, J.~P. Verbiest,
  and D.~G. Whelan, ``{A massive pulsar in a compact relativistic binary},''
  {\em Science}, vol.~340, p.~448, apr 2013.

\bibitem{Zhu2019}
W.~W. Zhu, G.~Desvignes, N.~Wex, R.~N. Caballero, D.~J. Champion, P.~B.
  Demorest, J.~A. Ellis, G.~H. Janssen, M.~Kramer, A.~Krieger, L.~Lentati,
  D.~J. Nice, S.~M. Ransom, I.~H. Stairs, B.~W. Stappers, J.~P. Verbiest,
  Z.~Arzoumanian, C.~G. Bassa, M.~Burgay, I.~Cognard, K.~Crowter, T.~Dolch,
  R.~D. Ferdman, E.~Fonseca, M.~E. Gonzalez, E.~Graikou, L.~Guillemot, J.~W.
  Hessels, A.~Jessner, G.~Jones, M.~L. Jones, C.~Jordan, R.~Karuppusamy, M.~T.
  Lam, K.~Lazaridis, P.~Lazarus, K.~J. Lee, L.~Levin, K.~Liu, A.~G. Lyne, J.~W.
  McKee, M.~A. McLaughlin, S.~Os{\l}owski, T.~Pennucci, D.~Perrodin,
  A.~Possenti, S.~Sanidas, G.~Shaifullah, R.~Smits, K.~Stovall, J.~Swiggum,
  G.~Theureau, and C.~Tiburzi, ``{Tests of gravitational symmetries with pulsar
  binary J1713+0747},'' {\em Monthly Notices of the Royal Astronomical
  Society}, vol.~482, no.~3, pp.~3249--3260, 2019.

\bibitem{2012MNRAS.423.3328F}
P.~C. Freire, N.~Wex, G.~Esposito-Far{\`{e}}se, J.~P. Verbiest, M.~Bailes,
  B.~A. Jacoby, M.~Kramer, I.~H. Stairs, J.~Antoniadis, and G.~H. Janssen,
  ``{The relativistic pulsar-white dwarf binary PSR J1738+0333 - II. The most
  stringent test of scalar-tensor gravity},'' {\em Monthly Notices of the Royal
  Astronomical Society}, vol.~423, pp.~3328--3343, jul 2012.

\bibitem{2000itss.book.....P}
R.~Ellis, {\em {An Introduction to the Theory of Stellar Structure and
  Evolution, 2nd edn., by Dina Prialnik}}, vol.~52.
\newblock jul 2011.

\end{thebibliography}

\documentclass[fleqn,12pt]{SelfArx} % Document font size and equations flushed left

\usepackage[english]{babel} % Specify a different language here - english by default
\usepackage{lipsum} % Required to insert dummy text. To be removed otherwise

\usepackage[utf8]{inputenc} % allow utf-8 input
\usepackage[T1]{fontenc}    % use 8-bit T1 fonts
\usepackage{hyperref}       % hyperlinks
\usepackage{url}            % simple URL typesetting
\usepackage{booktabs}       % professional-quality tables
\usepackage{amsfonts}       % blackboard math symbols
\usepackage{nicefrac}       % compact symbols for 1/2, etc.
\usepackage{microtype}      % microtypography
\usepackage{lipsum}	        % Can be removed after putting your text content
\usepackage[super,compress]{cite}
\usepackage{graphicx}% Include figure files
\usepackage{dcolumn}% Align table columns on decimal point
\usepackage{bm}% bold math
\usepackage{array}
\usepackage{tabularx}
\usepackage[english]{babel}
\usepackage[nottoc]{tocbibind}
\usepackage{amsmath}
\usepackage{threeparttable}
\usepackage{booktabs,caption}

\definecolor{color1}{RGB}{0,0,90} % Color of the article title and sections
\definecolor{color2}{RGB}{0,20,20} % Color of the boxes behind the abstract and headings

\usepackage{hyperref} % Required for hyperlinks
\hypersetup{hidelinks,colorlinks,breaklinks=true,urlcolor=color2,citecolor=color1,linkcolor=color1,bookmarksopen=false,pdftitle={Title},pdfauthor={Author}}

\JournalInfo{Arxive} % Journal information
\Archive{Preprint} % Additional notes (e.g. copyright, DOI, review/research article)

\PaperTitle{Period change of binary pulsars due to the accretion of dark matter particles} % Article title

\Authors{Ebrahim Hassani \textsuperscript{1}*, Hossein Ebadi \textsuperscript{2}, Reza Pazhouhesh \textsuperscript{3}} % Authors
\affiliation{\textsuperscript{1}\textit{Department of Physics, University of Birjand, Iran}} % Author affiliation
\affiliation{*\textbf{Corresponding author}: ebrahim.hassani@birjand.ac.ir}

\affiliation{\textsuperscript{2}\textit{Department of Theoretical Physics and Astrophysics, \\ University of Tabriz, PO Box 51664, Tabriz, Iran, e-mail: hosseinebadi@tabrizu.ac.ir}} % Author affiliation

\affiliation{\textsuperscript{3}\textit{Department of Physics, Faculty of Sciences, \\ University of Birjand, Birjand, Iran, e-mail: rpazhouhesh@birjand.ac.ir}} % Author affiliation

\Keywords{Dark Matter, Binary Pulsars, White dwarfs, Neutron stars} 
\newcommand{\keywordname}{Keywords} % Defines the keywords heading name

\Abstract{As binary systems move inside galaxies, they interact with the dark matter halo. This interaction leads to the accretion of dark matter particles inside binary components. The accretion of dark matter particles increases the mass of the binary components and then, the total mass of the system. Increased mass by this way can affect other physical parameters of the systems, like orbital periods of the systems. In this work, we estimated the period change of some known compact binary systems due to the accretion of dark matter particles inside them. Our conclustion is that, for WIMP particles with masses in the range $\simeq 10 \: GeV.c^{-2}$ and dark matter density as high as the dark matter density around the sun, period change due to accretion of dark matter partices is negligible compared to the period change due to the emission of gravitaional waves from the systems.}

\begin{document}

\flushbottom % Makes all text pages the same height

\maketitle
\tableofcontents
\thispagestyle{empty}

\section{Introduction} \label{Introduction} % The \section*{} command stops section numbering
Rotation curves of galaxies reveal the non-uniform distribution of dark matter (DM) inside galaxies \cite{Sofue2001}. From this point, one can infer that all stars that are distribited inside galaxies are immersed inside DM. Then, it is logical to suppose that the DM must affect structure and evolutionary courses of stars. For the first time Steigman used DM supposition to solve the solar neutrino problem \cite{1978AJ.....83.1050S} . Simulation of dwarf galaxies also supports the interaction between DM and stars \cite{Read_2018fxs} . In addition, DM effect on stars can be used to solve the paradox of youth problem for stars that are located near the Galactic massive black hole \cite{Hassani_2020zvz} . In addition to normal stars, DM effect on other celestiall bodies like the moon \cite{2020PhRvD.102b3024C, 2020PhLB..80435403G} , planets \cite{2012JCAP...07..046H} , neutron stars (NS) \cite{2012PhRvD..85b3519M, 2011IJMPE..20..109R, 2010PhRvL.105n1101P, 2010PhRvD..81l3521D, Bertone_2008, 2008PhRvD..77b3006K, Bell:2020jou, Joglekar:2020liw, 2020arXiv200405312L, 2019PhRvD.100j3019H, 2019JCAP...06..054B, 2020JCAP...06..007F, 2018JCAP...09..018B, 2018JHEP...11..096K, 2017PASA...34...43C, 2017arXiv170508864H, 2018ApJ...863..157C, 2017PhRvL.119m1801B, Perez-Garcia:2015rta, Perez-Garcia:2014nza, 2015ApJ...800..141Z, Tachibana:2013iva, 2013arXiv1306.0148P, Perez-Garcia:2013dwa, 2013PhRvD..87l3507B, 2013PhRvD..87e5012B, Serebrov:2012vm, 2012AIPC.1441..525P} , white dwarf stars (WD) \cite{Wong:2011ege, 2019JCAP...08..018D, 2018PhRvD..98f3002C, 2018PhRvD..98k5027G, 2018PhRvD..98j3023N, 2017arXiv170310104N, 2016MNRAS.459..695A, 2015PhRvD..91j3514H, 2013PhRvD..87l3506L}, black holes \cite{2020arXiv200405312L, 2012PhRvD..85b3519M, 2009JCAP...08..024U, Belotsky2014} and binary star systens \cite{Hassani2020} are also investigated in the literature. \\
For the first time, Press and Spergel estimated the rate at which DM particles (WIMP particles, exactly speaking) accrete into the sun \cite{1985ApJ...296..679P} . In the next step, Gould obtained a general relation for capture rate (CR) of DM particles by other round bodies like planets and stars \cite{Gould_1987ir} . Kouvaris used Press and Spergel relation to derive CR relation for NSs \cite{2008PhRvD..77b3006K} (see section \ref{Sec_CR_by_NS} for more details). Hurst et al used Gould relation to obtain CR relation for WDs \cite{2015PhRvD..91j3514H} (see section \ref{Sec_CR_bY_WDs} for more details).\\
As a compact binary system moves inside the DM halo, it feels dynamical friction by the DM halo. Induced dynamical friction by the DM halo can alter the orbital parameters (e.g. period and semi-major axis) of these systems. In a series of studies, reserchers tried to estimate the period change of pulsar binary systems due to this dynamical friction and compare the results with the period change due to the gravitational wave emission \cite{PhysRevD.96.063001, Penarrubia2016, Blas2020, GOMEZ2019100343, Blas2017, 2015PhRvD..92l3530P, Armaleo2020,  CAPUTO20181, Yoo2004}.\\
In this work we estimated the period change of compact binary systems due to the accretion of DM partilces (instead of due to the dynamical friction) inside them. Accretion of DM particles inside compact binary systems increases the mass of the binary components. Then, increased mass leads to the change in the orbital periods of these system. To the best of my knowledge, there is not any previous study that estimates the period change of binary systems due to the accretion of DM particles.\\
The rest of this paper is structured as follows. In sections \ref{Sec_CR_bY_WDs} and \ref{Sec_CR_by_NS} we presented CR ralations for WDs and NSs, separately. In section \ref{Sec_CR_By_Compact_Binary_Systems} we presented an approximate relation for the CR by compact binary systems. Finally, section \ref{Sec_Results_and_discussion} devoted to the results and discussions.

\section{Capture rate by white dwarfs} \label{Sec_CR_bY_WDs}
Capture Rate of DM particles by white dwarfs can be calculated using the equation \cite{2015PhRvD..91j3514H} :
\begin{equation} \label{Eq_CR_by_WDs}
CR_{WD} =  \underset{i}{\sum} CR_{WD, i} =\sqrt{\frac{3}{2}} \: \frac{\rho_{x}}{m_{x}} \: \sigma_{\chi, i} \: \frac{v_{esc,R_{WD}}}{\bar{v}} \: N_{i} \: \left \langle \widehat{\phi } \right \rangle \: \frac{erf(\eta)}{\eta }
\end{equation}
like equation \ref{Eq_CR_NS} for neutron stars, $ \rho_{\chi} $, $ m_{\chi} $ and $ \bar{v} $ are the density of DM halo around white dwarf (or around compact binary system), mass of WIPM particles and velocity dispersion of WIMPs (which we supposed to be $ \bar{v} = 220 \: km.sec^{-1}$ for all stars in this study), respectevely. $ R_{WD} $ is the radius of the white dwarf and $ v_{esc,R_{WD}} $ is the escape velocity at the surface of the white dwarf. $ \eta $ is the ratio of star's velocity through its halo to the local velocity dispersion of the DM halo. \\
$ \sigma_{\chi, i} $ is the scattering cross section of WIMP particles from nuclear species $i$ . As WDs consumed almost all of their hydrogen then, we just consider the spin-independent part of the scattering cross section and neglect the spin-dependent part. So, the spin-independent scattering cross section for all elements except hydrogen is \cite{1996PhR...267..195J, 2015PhRvD..91j3514H, Lopes2011} :
\begin{equation} \label{Eq_Scatterin_cross_wd}
\sigma_{\chi ,i} = \sigma_{\chi ,SI} A_{i}^{2}\left ( \frac{m_{\chi} m_{n, i}}{m_{\chi}+m_{n, i}} \right )^{2} \left ( \frac{m_{\chi}+m_{proton}}{m_{\chi}m_{proton}} \right )^{2}
\end{equation}
where $ \sigma_{\chi ,SI} $ is the spin-independent scattering cross-section which is determined through experimental direct DM detection methods. In this study we considered it to be $ \sigma_{\chi, SI} = 10^{-44} \: cm^{2} $ (as mentioned again in the section \ref{Sec_CR_by_NS} too). $ A_{i} $ is atomic number of species $i$, $ m_{n, i} $ in nuclear mass of species $i$ and $m_{proton}$ is the mass of a proton. To simplify calculations, we supposed that WDs composed of just carbon atoms with $m_{n,c} = 12 \: u$ where $u$ is the atomic mass unit (amu) and atomic number $A_{c} = 6$ . The similar supposition had done in the similar work \cite{2015PhRvD..91j3514H} . \\
In equation \ref{Eq_CR_by_WDs}, $ \widehat{\phi } $ is dimensionless potential for stellar nucleons and defined as \cite{2015PhRvD..91j3514H} :
\begin{equation}
\widehat{\phi } = \dfrac{v_{esc}^{2}(r)}{v_{esc}^{2}(R)}
\end{equation}
The quantity $\left \langle \widehat{\phi } \right \rangle$ is the average of $\widehat{\phi}$ over all necleons in the star. To calculate $\left \langle \widehat{\phi } \right \rangle$, we supposed that, like the case for NSs, average density of all WDs is constant. So, for a typical WD (e.g. Sirius B WD star with $M_{Sirius B} = 1.018 \: M_{\odot}$ mass \cite{Bond2017} and radius \cite{Holberg1998} $M_{Sirius B} = 0.0084 \: R_{\odot}$) we have:
\begin{equation} 
\bar{\rho}_{WD} = \dfrac{M_{WD}}{(4/3) \pi R^{3}_{WD}} = \dfrac{1.018 \: M_{\odot}}{(4/3) \pi (0.0084 \: R_{\odot})^{3}} = 2.38 \times 10^{9} \: \: \: (\dfrac{Kg}{m^{3}})
\end{equation}
So, the approximate mass-radius relation for white dwarfs becomes :
\begin{equation}
R_{WD} = \left(  \dfrac{M_{WD}}{(4/3) \pi \bar{\rho}_{WD}} \right) ^{\dfrac{1}{3}} = 4.65 \times 10^{-4} \times M_{WD}^{1/3} \: \: \: (m) .
\end{equation}
(We need $R_{WD}$ to calculate $v_{esc}(R)$ for WDs with different masses too). Now, $\left \langle \widehat{\phi } \right \rangle$ becomes:

\begin{equation}
\left \langle \widehat{\phi } \right \rangle = \dfrac{1}{R} \int_{0}^{R} \frac{v_{esc}^{2}(r)}{v_{esc}^{2}(R)}dr = \frac{1}{3}
\end{equation}
In equation \ref{Eq_CR_by_WDs}, $N_{i}$ is the total number of nucleons of species $i$ in the WD star. In the case of carbon WDs it becomes:
\begin{equation}
N_{c} = \dfrac{M_{WD}}{m_{c}} = \dfrac{M_{WD}}{12 \: u}
\end{equation}

\section{Capture rate by neutron stars} \label{Sec_CR_by_NS}
Capture Rate by Neutron stars, considering the general relativity corrections, can be calculated using the below equation \cite{2008PhRvD..77b3006K} :
\begin{equation} \label{Eq_CR_NS}
CR_{NS} = 4 \pi^{2} \: \frac{\rho_{x}}{m_x} \: (\frac{3}{2 \pi \bar{v}^{2}})^{3/2} \: \: \: \frac{(2GM_{NS}R_{NS})}{1-\frac{2GM_{NS}}{c^{2} \:R_{NS}}} \times min(\frac{1}{3} \: \bar{v}^{2} , E_{0}) \times f
\end{equation}

where $\rho_{x}$ is the density of the DM halo around the neutron star, $m_{x}$ is the mass of the WIMP particles, $ \bar{v} $ is the the velocity dispersion of the DM particles (in this study we took it to be $ \bar{v} = 220 \: km.sec^{-1}$ for all stars in the study), G is the gravitational constant, c is the speed of light, $M_{NS}$ is the mass of the neutron star and $R_{NS}$ is the radius of the neutron star. \\
To estimate the radius of a neutron stars $R_{NS}$ (to use in equation \ref{Eq_CR_NS}), we supposed that the average density of the all neutron stars is constant. So, the average density of a typical neutron star with $ M_{NS} = 1.44 \: M_{\odot} $ and radius $ R_{NS} = 10.6 \: km $ (see table 1 of the paper \cite{2014JCAP...05..013G} for these quantities) is :

\begin{equation} 
\bar{\rho}_{NS} = \dfrac{M_{NS}}{(4/3) \pi R^{3}_{NS}} = \dfrac{1.44 \: M_{\odot}}{(4/3) \pi (10.6 \times 10^{3})^{3}} = 5.74 \times 10^{17} \: \: \: (\dfrac{Kg}{m^{3}})
\end{equation}
So, we obtain an approximate mass-radius relation for neutron stars:
\begin{equation} 
R_{NS} = \left(  \dfrac{M_{NS}}{(4/3) \pi \bar{\rho}_{NS}} \right) ^{\dfrac{1}{3}} = 7.46 \times 10^{-7} \times M_{NS}^{1/3} \: \: \: (m) .
\end{equation}
In equation \ref{Eq_CR_NS}, $ E_{0} $ is a constant that parametrizes the maximum kinetic energy per mass of the WIMPs at an asymptotically large distance from the star in order for the WIMPS to be captured by the star. In the case of neutron stars and WIMPs with masses of the order GeV, $ E_{0} \gg (1/3) \bar{v}^{2}$. Then, in equation \ref{Eq_CR_NS}, $ (1/3) \bar{v}^{2} $ should be taken always as minimum (for more detailded dicussion about this selection see the section 2 of the paper \cite{2008PhRvD..77b3006K} .

Not all the WIPMs that enter a neutron star will be captured by it. In equation \ref{Eq_CR_NS}, $f$ represents the fraction of the WIPMs that, after one or several scattering from neclear matter, lose enough energy to be captured by a typical neutron star. For WIMP-necleon scattering cross sections $ \sigma_{\chi} > 10^{-45} \: cm^{2} $, $f = 1 $ (as in the case of this study) ; otherwise $f$ is given by \cite{2008PhRvD..77b3006K} :
\begin{equation} 
f = 0.45  \dfrac{\sigma_{\chi}}{\sigma_{crit}}
\end{equation}
which $ \sigma_{crit} = m_{n} R^{2} / M \simeq 6 \times 10^{-46} cm^{2} $. In this study we supposed $ \sigma_{\chi} = 10^{-44} cm^{2} $ (then, $f = 1$) which is the maximum magnitude that are determined through the experimental DM detection experiment \cite{2011PhRvL.106b1303B}.

\section{Capture rate by compact binary systems} \label{Sec_CR_By_Compact_Binary_Systems}
According to the Kepler's third law, square of the period in binary systems is proportional to the cube of the semi-major axis of the orbit \cite{hilditch_2001} :
\begin{equation} \label{Eq_Keplers_third_law}
P^{2} = (\dfrac{4 \pi^{2}}{GM}) \: a^{3} 
\end{equation}
which $M=M_{1}+M_{2}$ . If we suppose that all binary systems that are located inside the galaxies immeresed inside the DM halos then, it is logical to suppose that the both components of the systems accumulate DM particles while they are orbiting around each other. By this way, by passing the time, the total mass of the systems will increase. In the next, we will estimate how much the incresed mass will affect the other parameters of the systems. Taking differential of the both sides of the equation \ref{Eq_Keplers_third_law} yealds:
\begin{equation}
\begin{split}
2 P dP = \dfrac{4 \pi^{2}}{G} d(\dfrac{a^3}{M})  \\ 
= \dfrac{4 \pi^{2}}{G} \dfrac{3a^{2}da}{(M_{1}+M_{2})} - \dfrac{4 \pi^{2}a^{3}}{G} \dfrac{dM_{1}}{(M_{1}+M_{2})^{2}} - \dfrac{4 \pi^{2}a^{3}}{G} \dfrac{dM_{2}}{(M_{1}+M_{2})^{2}} \\
= \underbrace{\frac{4 \pi^{2} a^{3}}{G(M_{1}+M_{2})}}_{=P^{2}} \left [ \frac{3}{a}da - \frac{dM_{1}+dM_{2}}{M_{1}+M_{2}} \right ]
\end{split}
\end{equation}
which simplifies to :
\begin{equation} \label{Eq_Kepler_law}
\begin{split}
\dfrac{dP}{P} = \dfrac{3}{2} \dfrac{da}{a} - \dfrac{dM}{2M} = \dfrac{3}{2} \dfrac{da}{a} - \dfrac{dM_{1}+dM_{2}}{2(M_{1}+M_{2})}
\end{split}
\end{equation}
multiplying both sides of the equation \ref{Eq_Kepler_law} to $1/dt$ yealds :
\begin{equation} \label{Eq_CR_by_binary_systems}
\frac{\dot{P}}{P} = \frac{3}{2} \frac{\dot{a}}{a} - \frac{1}{2} \frac{\dot{M}}{M} = \frac{3}{2} \frac{\dot{a}}{a} - \frac{1}{2} \frac{\dot{M_{1}}+\dot{M_{2}}}{(M_{1} + M_{2})}
\end{equation}
where $\dot{P} = dp/dt$, $\dot{a} = da/dt$, $\dot{M} = dM/dt$, $\dot{M_{1}} = dM_{1}/dt$ and $\dot{M_{2}} = dM_{2}/dt$. $\dot{M_{1}}$ and $\dot{M_{2}}$ signifies the accretion of DM particles into compact stars. To calculate $\dot{M_{1}}$ and $\dot{M_{2}}$ it is enough to multiply equation \ref{Eq_CR_NS} (in the case of neutron stars) or equation \ref{Eq_CR_by_WDs} (in the case of white dwarf stars) to the mass of the DM particles $m_{\chi}$ (e.g. $\dot{M_{1}} = CR_{NS} \times m_{\chi}$ if the $M_{1}$ star is a neutron star). In table \ref{table_CR_in_known_binary_systems} we estimated  $\dot{M} / M $ (which $\dot{M} = \dot{M}_{1} + \dot{M}_{2}$) for some known compact binary systems. \\

\section{Results and Conclusions} \label{Sec_Results_and_discussion}
Using equation \ref{Eq_CR_by_binary_systems}, it is possible to estimate the accretion of DM particles in compact binary systems. 
\begin{table}
\centering
\caption{Compact binary system parameters that we used in this study to test our results. All parameters in the table are obtained through observations, except the $\dot{M}/M$ column which is obtained theoretically using the assumptions that are explained in the text.}
\begin{threeparttable}
\begin{tabular}{|| c | c || c | c || c | c | c | c | c | c ||}
\hline \hline
Name & \shortstack{ $M_{1}$ \\ $(M_{\odot})$ } & \shortstack{ $M_{2}$ \\ $(M_{\odot})$}  & \shortstack{ $ P_{d} $ \\ $(days)$ } & $\dot{P}$ \tnote{*} & \shortstack{ $\dot{P}/P$ \\ $(sec^{-1})$} & \shortstack{ $\dot{M}/M$ \\ $(sec^{-1})$} &  $Ref.$ \tnote{**} \\ \hline 

J1903+0327  & 1.0   & 1.7 & 95   & -6.4e-11 &  -7.8e-18  & 1.9e-32 & \cite{Freire2011} \\
J0737-3039  & 1.2   & 1.3 & 0.10 & -4.0e-15 &  -4.6e-19  & 3.5e-45 & \cite{Wex2014} \\
B1913+16    & 1.4   & 1.4 & 0.32 & 5.0e-15  &  1.8e-19   & 3.6e-45 & \cite{Wex2014} \\
J1012+5307  & 0.16  & 1.6 & 0.60 & -1.8e-14 &  -3.5e-19  & 2.5e-32 & \cite{Arzoumanian2018, Desvignes2016} \\
J1614-2230  & 0.49  & 1.9 & 8.7  & 3.4e-13  &  4.5e-19   & 2.6e-32 & \cite{Arzoumanian2018} \\
J1909-3744  & 0.21  & 1.5 & 1.5  & -4.0e-15 &  -3.1e-20  & 2.3e-32 & \cite{Arzoumanian2018, Desvignes2016} \\
J0751+1807  & 0.16  & 1.6 & 0.26 & -4.6e-14 &  -2.1e-18  & 2.5e-32 & \cite{Desvignes2016} \\
J1910+1256  & 0.19  & 1.6 & 58   & -2.0e-11 &  -4.0e-18  & 2.5e-32 & \cite{Gonzalez2011} \\
J2016+1948  & 0.29  & 1.0 & 635  & -1.0e-9  &  -1.8e-17  & 2.7e-45 & \cite{Gonzalez2011} \\
J0348+0432  & 0.17  & 2.0 & 0.10 & -1.1e-14 &  -1.3e-18  & 3.2e-32 & \cite{2013Sci...340..448A} \\
J1713+0747  & 0.29  & 1.3 & 68   & 3.0e-14  &  5.1e-21   & 2.9e-45 & \cite{Zhu2019} \\
J0613-0200  & 0.12  & 1.2 & 1.2  & 2.7e-14  &  2.6e-19   & 2.5e-45 & \cite{Arzoumanian2018, Desvignes2016} \\
J1738+0333  & 0.19  & 1.5 & 0.35 & 2.0e-15  &  6.6e-20   & 2.4e-32 & \cite{2012MNRAS.423.3328F} \\ \hline

\end{tabular}
\begin{tablenotes}
\item[*] $\dot{P}$ is a dimensionless quantity.
\item[**] References for all parameters except $\dot{M}/M$, which is calculated in this study.
\end{tablenotes}
\end{threeparttable}
\label{table_CR_in_known_binary_systems}
\end{table}
In our calculations, we supposed stars with masses lower than the chandrasekhar mass ($1.4 \: M_{\odot}$) to be white dwarfs \cite{2000itss.book.....P} . Stars with masses higher than $1.4 \: M_{\odot}$ considered to be neutron stars. The mass of the WIMP particles and density of DM halo around binary systems considered to be $m_{\chi}=10 \: \: GeV.c^{-2}$ and $\rho_{\chi} = 8.16 \times 10^{-3} M_{\odot}/pc^{3}$ (which is the estimated DM density around solar neibourhoud \cite{Bertone2005a} ), respectively. \\
In equation \ref{Eq_CR_by_binary_systems}, it is possible to measure $\dot{P}/P$ term through observations.  But the precision of the current observational technology is not enough to measure the $ \dot{a}/a $ term in the right hand side of the equation. \\
Paying attention to the amounts of $\dot{M}/M$ and $\dot{P}/P$ in table \ref{table_CR_in_known_binary_systems}, we see that the amounts of $\dot{M}/M$ are about 12-14 orders of magnitude smaller than the amounts of $\dot{P}/P$ (for binaries that at least one of the components is a neutron star). For binaries that both of the components are white dwarfs (e.g. J0737-3039, B1913+16 or J2016+1948 systems) the results are more dramatic and the amounts of $\dot{M}/M$ are about 24-26 orders of magnitude smaller than the amounts of $\dot{P}/P$. Then, we can say that the accretion of DM particles is not the main reason of period change in compact binaries. In order to interpret the period changes $\dot{P}/P$, other physical sources that affects systems parameters, e.g. gravitaional emission, mass transfer, dynamical friction etc., must be token into account. \\
As an overall result, our calculations show that the accretion of DM particles in compact binary systems do not change the systems parameters significantly. In addition, accretion of DM particles can not be considered the main source of period changes in binary pulsars.

\phantomsection
\section*{Acknowledgments} % The \section*{} command stops section numbering
\addcontentsline{toc}{section}{Acknowledgments} % Adds this section to the table of contents
The author would like to express his special thanks to Prof. Joakim Edsjö from the University of Stockholm, Sweden, and Dr. Amin Rezaei Akbarieh from the University of Tabriz, Iran and Prof. Gianfranco Bertone from the University of Amsterdam, Netherlands and Dr. Marco Taoso from the Istituto Nazionale di Fisica Nucleare (INFN), Italy because of their worthwhile discussions during the research.
\phantomsection
\bibliographystyle{ieeetr}
\bibliography{My_References}
\end{document}